# Grouped Sequency-arranged Rotation: Optimizing Rotation Transformation for Quantization for Free

### v1

@@ 2.1 Walsh-Hadamard Transform and Sequency
[…]
@@ 2.2 Rotation for LLM Quantization
Since a Hadamard matrix can be used as a rotation matrix when scaled and has an efficient algorithm, recent state-of-the-art methods make extensive use of the Hadamard transform Ashkboos et al. (2024); Xiang et al. (2025); Lin et al. (2024); Liu et al. (2024); Hu et al. (2025). Following SpinQuant’s terminology, $R_{1}$ rotates all hidden activations between transformer blocks, $R_{2}$ rotates the value activation, $R_{3}$ rotates the query and key activations after RoPE, and $R_{4}$ rotates the input activation of the down projection. Specifically for $R_{1}$, a Randomized Hadamard Transform (RHT) is employed following the proposition in Quip# Tseng et al. (2024) for better incoherence processing. This way, the outliers in the activation distribution are largely suppressed, achieving deployable W4A4KV4111We notate x-bit weight, y-bit activation, z-bit KV-cache into WxAyKVz like W4A4KV4. performance on famous LLM models.
@@ 3.1 Grouped Sequency-arranged Rotation
, where $H_{wal}\in\{-1,1\}^{G\times G}$ is a ${G\times G}$ Walsh matrix, with $G$ being the quantization group size, and 0 is the $G\times G$ zero matrix.
The proposed $R_{GSR}$ has several advantages over the RHT and the SpinQuant matrices: First, it can replace any rotation matrix in existing PTQ methods without training like QuaRot Ashkboos et al. (2024) for virtually free, as the only additional operation required is to pre-process a Sylvester-constructed Hadamard matrix to a Walsh matrix and apply the Kronecker product with an identity matrix before running quantization. Second, it can systematically reduce weight quantization error by strategically arranging sequency filters with similar yet diverse sequency values (Section 3.2). Third, it can also serve as an enhanced initialization for training-based methods such as SpinQuant Liu et al. (2024) and OSTQuant Hu et al. (2025) (Section 4).
@@ 3.2 The Effect of Sequency Arrangement on Group Quantization
where $R_{f}$ and $R_{r}$ are rotation matrices applied to the front and rear side of a weight $W$, respectively. For query weight $W_{q}$ as an example, $R_{f}=R_{1}$ and $R_{r}=I$ hold. We do not consider local rotation for brevity.
An $(i,j)$ element of the rotated weight ($W^{\prime}[i,j]$) can be derived as follows:
@@
$=$ $\left\langle\left[\langle R_{f}^{-1}[i,:],W[:,1]\rangle,\langle R_{f}^{-1}[i,:],W[:,2]\rangle,\right.\right.$
@@
An $n$-th group in $W^{\prime}$ can be expressed as $W^{\prime}[nG:(n+1)G,:]$, which leads to our observation #1 by simply substituting $i$ to $nG:(n+1)G$ in Eqn. 5.
Observation #1 Under group quantization, each column group in the front rotation matrix $R_{f}$ generates distinct rotated weight groups, and the rear rotation matrix $R_{r}$ always applies on the whole weight. In other words, a group in the rotated weight $W^{\prime}$ is the original weight transformed by the corresponding group of filters in the front rotation matrix and then by all filters in the rear rotation matrix.
@@
Now, we relate the sequency arrangement to group quantization performance. For $R_{r}$, the arrangement has no impact as long as the set of sequency values is equal, which is the case with comparing the Hadamard and Walsh matrices. Therefore, we focus on $R_{f}$. The Walsh matrix (with the sequency ordering) has smaller sequency variance within each column group than the Hadamard matrix. Since sequency is analogous to frequency in the conventional frequency-domain filtering, the Walsh matrix will produce rotated weight groups with fewer massive outliers. Since $R_{1}$ works as $R_{f}$ on transformer weights such as $W_{q},W_{k},W_{v},W_{up},\ \text{and}\ W_{gate}$, changing $R_{1}$ from Hadamard to Walsh helps reduce the quantization error for these weights.
@@
The randomization method in Quip# Tseng et al. (2024) and QuaRot Ashkboos et al. (2024) only flips the signs of diagonal elements in a Hadamard matrix. This process keeps the overall sequency arrangement with no significant changes. Therefore, we can compare the RHT against the Walsh using the same logic as in the previous section.
@@
Local rotation (using block-diagonal matrices) is generally more effective than global rotation (using a single large matrix) Lin et al. (2024); Xiang et al. (2025); Xiang and Zhang (2024). Global rotation can struggle to handle outliers effectively, whether in activations or weights, as it tends to spread their impact widely. Local rotation, however, confines the effects of such outliers within their specific block or group as in Fig. 2 (b). When used with the Walsh matrix, this containment helps better reduce errors, which is also beneficial for low bit weight quantization.
4 Experimental Results
@@
The overall results are summarized in Table 1. Across all methods, our proposed approach consistently outperforms the GH, achieving lower PPL and higher accuracy on reasoning tasks. In particular, applying the GW to QuaRot (i.e., re-ordering rows of the Hadamard matrix with natural ordering) yields approximately 1 point lower PPL compared to SpinQuant, validating the benefit of the sequency arrangement. Given that SpinQuant typically consumes much greater computational costs than QuaRot, this result suggests that adopting GSR enables QuaRot to achieve superior performance and efficiency.
@@ 5 Conclusion
Our proposed method has proven effective only under extremely low bit weight quantization. On larger bit settings, the quantization error becomes much less significant, so that the sequency alignment cannot show visible improvement.
@@ A.3 Complete Reasoning Tasks Results
Table 3: Complete comparison of accuracy on zero-shot common-sense reasoning tasks for Llama-2-7B with QuaRot and SpinQuant. lambada-o and lambada-s represent lambada-openai and lambada-standard, respectively. LR and LS represent learning rotation and learning smooth factors, respectively.
@@
Table 4: Complete comparison of accuracy on zero-shot common-sense reasoning tasks for Llama-2-7B with OSTQuant. lambada-o represent lambada-openai. LR and LS represent learning rotation and learning smooth factors, respectively.

### v2

@@ 2.1 Walsh-Hadamard Transform and Sequency
$\text{S}(\text{i})=\text{bit\_count}(\text{i}\oplus(\text{i}>>1)).$ (2)
@@ 2.2 Rotation for LLM Quantization
Since a Hadamard matrix can be used as a rotation matrix when scaled and has an efficient algorithm, recent state-of-the-art methods make extensive use of the Hadamard transform Ashkboos et al. (2024); Xiang et al. (2025); Lin et al. (2024); Liu et al. (2025); Hu et al. (2025). We followed SpinQuant’s terminology to describe our rotation scheme as Fig. 1. At Fig. 1, $R_{1}$ rotates all hidden activations between transformer blocks, $R_{2}$ rotates the value activation, $R_{3}$ rotates the query and key activations after RoPE, and $R_{4}$ rotates the input activation of the down projection. Specifically for $R_{1}$, a Randomized Hadamard Transform (RHT) is employed following the proposition in Quip# Tseng et al. (2024) for better incoherence processing. This way, the outliers in the activation distribution are largely suppressed, achieving deployable W4A4KV4111We notate x-bit weight, y-bit activation, z-bit KV-cache into WxAyKVz like W4A4KV4. performance on famous LLM models.
@@ 3.1 Grouped Sequency-arranged Rotation
, where $H_{wal}\in\{-1,1\}^{G\times G}$ is a $G\times G$} Walsh matrix, with $G$ being the quantization group size, and 0 is a $G\times G$ zero matrix.
The proposed $R_{GSR}$ has several advantages over the RHT and the SpinQuant matrices: First, like QuaRot Ashkboos et al. (2024), it can replace any rotation matrix in existing PTQ methods without training for virtually free, as the only additional operation required is to pre-process a Sylvester-constructed Hadamard matrix to a Walsh matrix and apply the Kronecker product with an identity matrix before going into quantization. Second, it can systematically reduce weight quantization error by strategically arranging sequency filters with similar yet diverse sequency values (Section 3.2). Third, it can also serve as an enhanced initialization for training-based methods such as SpinQuant Liu et al. (2025) and OSTQuant Hu et al. (2025) (Section 4).
@@ 3.2 The Effect of Sequency Arrangement on Group Quantization
where $R_{f}$ and $R_{r}$ are rotation matrices applied to the front and rear side of a weight $W$, respectively. For query weight $W_{q}$ as an example, $R_{f}=R_{1}$ and $R_{r}=I$ hold. We do not consider local rotation in this section for brevity.
An $(i,j)$ element of the rotated weight ($W^{\prime}[i,j]$) can be expressed as follows:
@@
[…]
@@
An $n$-th row group in $W^{\prime}$ can be expressed as $W^{\prime}[nG:(n+1)G,:]$, which leads to our observation #1 by simply substituting $i$ to $nG:(n+1)G$ in Eqn. 5.
Observation #1 Under group quantization, each column group in the front rotation matrix $R_{f}$ generates distinct rotated weight groups, and all columns in the rear rotation matrix $R_{r}$ are always applied to all rows in the original weight. In other words, a group in the rotated weight $W^{\prime}$ is the original weight transformed by the corresponding group of filters in the front rotation matrix and then by all filters in the rear rotation matrix.
@@
Now, we relate the sequency arrangement to group quantization performance. For $R_{r}$, the arrangement has no impact as long as the set of sequency values is equal, which is the case with comparing the Hadamard and Walsh matrices. Therefore, we focus on $R_{f}$. The Walsh matrix (with the sequency ordering) has smaller sequency variance within each column group than the Hadamard matrix because the sequency values increase linearly. Since sequency is analogous to frequency in the conventional frequency-domain filtering, the Walsh matrix will produce rotated weight groups with fewer massive outliers. As shown in Table 1, $R_{1}$ works as $R_{f}$ on many different types of transformer weights including $W_{q},W_{k},W_{v},W_{up},\ \text{and}\ W_{gate}$, changing $R_{1}$ from Hadamard to Walsh helps reduce the quantization error for these weights.
Weight $W_{q}$ $W_{k}$ $W_{v}$ $W_{o}$ $W_{up}$ $W_{gate}$ $W_{down}$
$R_{f}$ $R_{1}$ $R_{1}$ $R_{1}$ $R_{2}$ $R_{1}$ $R_{1}$ $R_{4}$
$R_{r}$ II II $R_{2}$ $R_{1}$ II II $R_{1}$
Table 1: Rotation matrix configuration on each weight type in LLaMA-like transformer architecture. II is the identity matrix.
@@
The randomization method in Quip# Tseng et al. (2024) and QuaRot Ashkboos et al. (2024) only flips the signs of diagonal elements in a Hadamard matrix. This process keeps the overall sequency arrangement with no significant changes. Therefore, we can compare the RHT against the Walsh following the same logic as in the previous section.
@@
Local rotation (using block-diagonal matrices) is generally more effective than global rotation (using a single large matrix) Lin et al. (2024); Xiang et al. (2025); Xiang and Zhang (2024). Global rotation can struggle to effectively handle outliers, whether in activations or weights, as it spreads their impact to the whole input channel. Local rotation, however, confines the effects of such outliers within their specific block or group as in Fig. 2 (b). When used with the Walsh matrix, this containment helps better reduce errors, which is also beneficial for low-bit weight quantization.
@@ 4 Experimental Results
The overall results are summarized in Table 2. Across all methods, our proposed approach consistently outperforms the GH, achieving lower PPL and higher accuracy on reasoning tasks. In particular, applying the GW to QuaRot (i.e., re-ordering rows of the Hadamard matrix with natural ordering) yields approximately 1 point lower PPL compared to SpinQuant, validating the benefit of the sequency arrangement. Given that SpinQuant typically consumes much greater computational costs than QuaRot, this result suggests that adopting GSR enables QuaRot to achieve superior performance and efficiency. While OSTQuant learns both the rotation matrix and the smooth factor through optimization and achieves a PPL of 10.97 in the W2 setting, QuaRot with GSR attains a comparable PPL of 11.59 by simply replacing $R_{1}$in a training-free manner. In the W2A4 setting, QuaRot with GSR even surpasses OSTQuant, achieving a lower PPL of 15.23 compared to 16.16, indicating that better performance can be obtained with fewer resources. The effectiveness of GSR also holds when applied to OSTQuant, consistently leading to further performance gains.
@@ 5 Conclusion
Our proposed method has proven effective only under extremely low-bit weight quantization with group quantization. On larger bit configurations, the quantization error becomes much less significant, so that the sequency alignment cannot show visible improvement. In addition, to ensure the generalizability of our approach, we plan to extend our experiments to other model architectures and datasets in future work.
Acknowledgments
This work was supported by Samsung Advanced Institute of Technology, and MX division, Samsung Electronics Co., Ltd., Inter-university Semiconductor Research Center (ISRC) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].
@@ A.3 Complete Reasoning Tasks Results
In this section, Table 4 and Table 5 present evaluation results for each zero-shot task.
@@
Table 4: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks for Llama2-7B with QuaRot and SpinQuant. lambada-o and lambada-s represent lambada-openai and lambada-standard, respectively.
@@
Table 5: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks for Llama2-7B with OSTQuant. lambada-o represents lambada-openai.